A man is marooned at rest on level frictionless ice. In desperation, he hurls his shoe to the right at 4.1m/s. If the man weighs 385 N and the shoe weighs 2.1N, the man moves to the left with a speed of...


1
Expert's answer
2021-12-02T10:05:26-0500
"(385-2.1)V=(2.1)(4.1)\\V=0.0225\frac{m}{s}"